OpenStudy (anonymous):

the 3 numbers 1428, 2134, and 2850 has the same remainder y when they all divided by x. the possible value of x + y = ....

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ahsome):

Is this what the values mean: \[y=\frac{1428}{x}\]\[y=\frac{2134}{x}\]\[y=\frac{2850 }{x}\]

OpenStudy (anonymous):

px + y = 1428 qx + y = 2134 rx + y = 2850

OpenStudy (anonymous):

In other words, \[1428\equiv2134\equiv2850\equiv x\mod y\,?\]

OpenStudy (anonymous):

Sorry, should be \(\color{red}{y\mod x}\).

OpenStudy (anonymous):

y its remainder for that numbers

OpenStudy (anonymous):

yes, y modx what's next ?

OpenStudy (anonymous):

Notice that \[\begin{cases}2850-2134\equiv y-y\mod x~~\implies~~716\equiv 0\mod x\\ 2134-1428\equiv y-y\mod x~~\implies~~706\equiv0\mod x \end{cases}\] In turn, \[716-706\equiv0-0\mod x~~\implies~~10\equiv0\mod x\]

ganeshie8 (ganeshie8):

is the answer 2 ?

OpenStudy (anonymous):

still didnt understand, what do you means ?

OpenStudy (anonymous):

here the choices answer : a. 165 b. 179 c. 344 d. 716

OpenStudy (anonymous):

You can divide 10 by four different integers such that the remainder is 0 (i.e. \(10\equiv0\mod x\)), which suggests \(x=1\), \(2\), \(5\), or \(10\). Of these four, only \(x=2\) agrees with the fact that \(706\equiv0\mod 2\), and so on with the others. The fact that the three given numbers are all even suggests that \(y=0\).

OpenStudy (anonymous):

I'm agreeing with @ganeshie8 but there may be more to this modular arithmetic than what I can gather from glancing at the wiki page on it.

OpenStudy (anonymous):

ah, sorry... extra information says with y not equal zero

ganeshie8 (ganeshie8):

Haha! now it is interesting

OpenStudy (anonymous):

i think if y = 0, the question not enough interesting :p

ganeshie8 (ganeshie8):

yes y = 0, x = 2 is a trivial solution.. going for dinner.. il give a good try after coming back if it still remains unsolved

OpenStudy (anonymous):

ok, thanks :D

ganeshie8 (ganeshie8):

@mathmath333 try this :)

OpenStudy (mathmath333):

\(\large \color{black}{ 1428=px+y\hspace{.33em}\\~\\ 2134=qx+y\hspace{.33em}\\~\\ 2850=rx+y\hspace{.33em}\\~\\\\~\\\\~\\ \normalsize \text{substracting those}\hspace{.33em}\\~\\\\~\\\\~\\ 706=(q-p)x\hspace{.33em}\\~\\ 716=(r-q)x\hspace{.33em}\\~\\ 1422=(r-p)x\hspace{.33em}\\~\\~\\~\\ \color{red}{2}\times 353=(q-p)\color{red}{x}\hspace{.33em}\\~\\ \color{red}{2}\times2\times 179=(r-q)\color{red}{x}\hspace{.33em}\\~\\ \color{red}{2}\times 3^2\times79 =(r-p)\color{red}{x}\hspace{.33em}\\~\\\\~\\\\~\\ \normalsize \text{as gcd of those is 2,only x=2 makes sense}\hspace{.33em}\\~\\ \normalsize \text{which is the largest number which leaves the same remainder}\hspace{.33em}\\~\\ \normalsize \text{but when x=2 ,y=0 , and u dont want that condition.}\hspace{.33em}\\~\\ \normalsize \text{so i think may be something is wrong with that numbers may be a typo}\hspace{.33em}\\~\\ }\)

OpenStudy (anonymous):

I whipped up a script in Mathematica to check the remainders for values of \(x\) up to \(100,000\) (unnecessarily high, I know) that suggests \(x=1\) or \(x=2\), with \(y=0\) in both cases. I agree with @mathmath333, there must be a problem with the given answers or the question.